OpenStudy (anonymous):

Find an equation of the tangent line to y = f(x) at x = a. f(x) = sqrt(x^2 + 16), x = 13

OpenStudy (anonymous):

\[f(x)=\sqrt{x^2+16}\Rightarrow f(13)=\sqrt{185}\\ f'(x)=\frac{2x}{\sqrt{x^2+16}}\Rightarrow f'(13)=\frac{26}{\sqrt{185}}\] The tangent line is given by the point-slope formula using the point (13, √185) and the slope 26/√185: \[y-\sqrt{185}=\frac{26}{\sqrt{185}}(x-13)\]

OpenStudy (anonymous):

so thats the equation? @SithAndGiggles

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

@myko

OpenStudy (anonymous):

the derivative is wrong. The nominator should be just x

OpenStudy (anonymous):

okay u mean in 2x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x/sqrt(x^2+16)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so it wouldn't be 26

OpenStudy (anonymous):

rest looks right, just change 26 for 13 and that's it

OpenStudy (anonymous):

okay thx

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Thanks for catching the mistake